# Towards Prospective Medical Image Reconstruction via Knowledge-Informed Dynamic Optimal Transport

### v1

@@ A.1 KIDOT Training Algorithm
5:        Evolve $\{y^{(j)}\}$ to $\{\mathcal{T}_{\phi}^{t_{N-1}}(y^{(j)})\}$ using discrete transport equation for $N$ steps.
@@
By [50, Theorem 5.11], narrow convergence and convergence of first moments together imply that
@@
=(16)(17)${=}\lim_{n\to\infty}\mathbb{E}_{\mathbb{P}}\Bigl{[}\int_{0}^{1}c\bigl{(}y,\mathcal{T}^{t}_{\phi_{n}}(y)\bigr{)}\mathrm{d}t\Bigr{]}+\lambda W_{1}\bigl{(}\mathbb{Q},(\mathcal{T}^{t}_{\phi_{n}})_{\#}\mathbb{P}\bigr{)}$ )
@@ B.1 Implementation Details and Baselines
We trained distinct KIDOT models tailored to each specific datasets (simulated MRI data, real prospective MRI data and clinical prospective CT data). Optimization utilized the RMSProp algorithm with differential learning rates: $1\times 10^{-4}$ for the transport network ($\mathcal{T}_{\phi}$) and $2\times 10^{-4}$ for the potential network ($\varphi_{\theta}$). The inner iteration parameter $N_{c}$ was consistently set to 1. A learning rate decay schedule applied a factor of 10 reduction after each block of 30 training epochs. The underlying architecture for $\mathcal{T}_{\phi}$ was adapted based on the dataset: using the E2E-VarNet backbone [12] for simulated MRI data, the PromptMR backbone [13] for real prospective MRI data, and the MC-CDic backbone [47] for clinical prospective CT data. Experiments were performed using PyTorch on an NVIDIA 4090 GPU. The source code will be publicly released following possible publication.
@@
In our MRI reconstruction experiments, we conducted comparisons against several existing methods, including the traditional CS-wavelet [43], two representative deep learning unfolding networks, E2E-VarNet [12] and PromptMR [13], and two comparative methods based on OT: OT-CycleGAN [34] and UAR [36].
For CT reconstruction, our set of comparison methods consists of the traditional BM3D [45], two representative deep learning methods (ISTA-Net [47], an unfolding network, and REDCNN [46]), and two OT-based approaches, OT-CycleGAN [34] and UAR [36].
@@ B.2 Datasets Details
All images were center-cropped to 320×320 pixels. We simulated undersampled measurements by applying k-space masks equivalent to a 4$\times$ acceleration factor. A key aspect of this evaluation was to assess KIDOT’s robustness to scenarios where test data characteristics differ from training data, a common challenge in prospective clinical deployment. To achieve this, we generated two distinct sets of undersampled inputs: one set, with specific sampling patterns, was used for training supervised baselines, while a different set of sampling patterns was employed for testing to mimic prospective acquisition conditions. Specifically, the test mask was generated by randomly flipping 3% of the entries in the training mask. Examples of these distinct masks are shown in Figure. 5.
@@
Given the potential for anatomical misalignment in real prospective acquisitions (due to different scan directions/timings), we also generated perfectly-paired Simulated Low-Dose CT (SLDCT) images by adding noise in the projection domain. We inject compound Gaussian and Poisson noise into the uncorrupted projections by using the method in [48, 49]. The incident scale photon flux is 40000.We adopt a fan-beam geometry to simulate the uncorrupted projections from the CT images. The geometrical parameters for projection were: source-to-isocenter distance 570.0 mm, source-to-detector distance 962.9 mm, image size 512 × 512 (with pixel spacing 0.6934 × 0.6934 $\text{mm}^{2}$), detector bin number 864, detector bin width 1.0336 mm, and 1200 projection views acquired over a 360-degree orbit.
@@ B.3 Additional Ablation Studies
B.3.1 Sensitivity to the Supervised Trade-off Parameter $\gamma$
We analyze the sensitivity of KIDOT to the supervised trade-off parameter $\gamma$ using simulated MRI data (4$\times$ acceleration). As shown in Table 5, the choice of $\gamma$ impacts KIDOT’s performance metrics. We observe a general trend where increasing $\gamma$ from 1 leads to consistent improvements in reconstruction quality (higher PSNR and SSIM) and fidelity (lower FID and KID). The best results across all evaluated metrics are achieved when $\gamma=10^{4}$. Further increasing the trade-off parameter to $\gamma=10^{5}$ results in a slight degradation across all metrics compared to $\gamma=10^{4}$. Based on this result, we fix $\gamma=10^{4}$ in our experiments.
Table 5: Comparison of different trade-off parameter $\gamma$ on the Simulated MRI data (4$\times$ acceleration). Best results are bolded.
trade-off parameter $\gamma$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ KID×$\times$100 $\downarrow$
@@ B.3.1 Sensitivity to the Supervised Trade-off Parameter $\gamma$
B.3.2 Effect of Different Lagrange Multiplier $\lambda$
We investigate the effect of the Lagrange multiplier $\lambda$ on KIDOT’s performance using simulated MRI data (4$\times$ acceleration). Table 6 summarizes the results for different values of $\lambda$. We observe that the choice of $\lambda$ has an impact on both reconstruction quality and fidelity metrics. Notably, the optimal performance across all evaluated metrics (highest PSNR/SSIM and lowest FID/KID) is achieved with a Lagrange multiplier of $\lambda=1$. Increasing $\lambda$ beyond this value leads to a consistent and substantial degradation in performance. As $\lambda$ increases to $10^{2}$, PSNR and SSIM decrease, while FID and KID×$\times 100$ values increase drastically, indicating a considerable loss of image fidelity and introduces more artifacts. Based on these results, we fix $\lambda=1$ for our experiments.
Table 6: Comparison of different Lagrange multiplier $\lambda$ on the Simulated MRI data. (4$\times$ acceleration). Best results are bolded.
Lagrange multiplier $\lambda$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ KID×$\times$100 $\downarrow$
@@ B.3.2 Effect of Different Lagrange Multiplier $\lambda$
B.3.3 Effect of the Physical Forward Operator $\mathcal{A}$ in Transport Equation
We conducted an ablation study to evaluate the importance of explicitly incorporating the physical forward operator $\mathcal{A}$ within our KIDOT framework. This comparison was performed on the simulated MRI data with a 4$\times$ acceleration factor. The setting "w/o $\mathcal{A}$" indicates that the physical forward operator was removed from the transport equation, or the network architecture $\mathcal{T}_{\phi}$. As summarized in Table 7, including the physical forward operator $\mathcal{A}$ leads to a substantial improvement in reconstruction quality and fidelity. Comparing the setting "w $\mathcal{A}$" to "w/o $\mathcal{A}$", we observe that incorporating $\mathcal{A}$ results in a notable increase in PSNR (33.31 vs 30.88) and SSIM (0.8518 vs 0.8174). The fidelity metrics are improved, with FID decreasing from 36.53 to 10.68, and KID×$\times 100$ decreasing from 1.48 to 0.83. These results clearly demonstrate the crucial role of the physical forward operator in guiding the reconstruction process and highlight its contribution to KIDOT’s superior performance.
Table 7: Results of the method w/wo physical forward operator $\mathcal{A}$ on the Simulated MRI data (4$\times$ acceleration). Best results are bolded.
Setting PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ KID×$\times$100 $\downarrow$
@@ B.4 Statistical Significance Analysis
We perform statistical significance tests to compare baseline methods against our proposed KIDOT method on the simulated MRI dataset with 4$\times$ acceleration. Table 8 summarizes the p-values from these tests. Each test compares the results of a baseline method directly against KIDOT.
@@
For distribution-based fidelity metrics (FID and KID×$\times 100$), which are traditionally computed once per dataset, we obtain multiple samples for statistical testing using a bootstrap resampling procedure over the test images within each run. The FID and KID scores computed on these bootstrap samples form the basis for independent two-sample t-tests comparing baseline methods with KIDOT.
As shown in Table 8, all p-values fall below the conventional significance threshold of 0.05. Notably, comparisons against CS-wavelet, E2E-VarNet, and OT-CycleGAN yield p-values $<0.001$ across all metrics, indicating strong statistical significance. For UAR, p-values range from 0.009 (KID×$\times 100$) to 0.024 (PSNR), still demonstrating significant improvements by KIDOT. These results collectively confirm that KIDOT outperforms all tested baselines on the simulated MRI dataset with high statistical confidence.
Table 8: Statistical significance (p-values) of baseline methods compared to KIDOT on the Simulated MRI data (4$\times$ acceleration). Lower p-values indicate more significant differences. P-values are from independent t-tests.
Method PSNR SSIM FID KID×$\times$100
CS-wavelet $<0.001$ $<0.001$ $<0.001$ $<0.001$
E2E-VarNet $<0.001$ $<0.001$ $<0.001$ $<0.001$
OT-CycleGAN $<0.001$ $<0.001$ $<0.001$ $<0.001$
UAR $0.024$ $<0.001$ $<0.001$ $0.009$
B.5 Training Cost Curves for Two Datasets
In Figure 7, we display the cost curves over two datasets (i.e., simulated MRI data, and clinical prospective CT data) of $\mathcal{T}_{\phi}$ and $\varphi_{\theta}$ in the training process. The $\mathcal{T}_{\phi}$ cost curve is normalized to [0, 1]. $\varphi_{\theta}$ cost is scaled to [0, 1] and then take the negative.
@@ B.5 Training Cost Curves for Two Datasets
B.6 Additional Visual Results
B.6.1 Additional Visual Comparison on Simulated MRI Data
@@
B.6.2 Additional Visual Comparison on Real Prospective MRI data
@@
B.6.3 Additional Visual Comparison on Clinical Prospective CT Data

### v3

@@ 5 Conclusion and Limitations
Acknowledgement
This work was supported by National Key R&\&D Program(2022YFA1004201), NSFC (12090021, 12125104, 12401671, 12426313, 12501708, 12501709, 623B2084), Postdoctoral Research Funding Project of Shaanxi Province(2024BSHTDZZ007), China National Postdoctotal Program for Innovative Talents (BX20240276), China Fundamental Research Funds for the Central Universities (xzy022025047), China Postdoctoral Science Foundation (2025M773058), and the Key Laboratory of Biomedical Imaging Science and System, Chinese Academy of Sciences.
@@ A.1 KIDOT Training Algorithm
5:   Evolve $\{y^{(j)}\}$ to $\{\mathcal{T}_{\phi}^{t_{N-1}}(y^{(j)})\}$ using discrete transport equation for NN steps.
@@
By [santambrogio2015optimal, Theorem 5.11], narrow convergence and convergence of first moments together imply that
@@
[…]
@@ B.1 Implementation Details and Baselines
We trained distinct KIDOT models tailored to each specific datasets (simulated MRI data, real prospective MRI data and clinical prospective CT data). Optimization utilized the RMSProp algorithm with differential learning rates: $1\times 10^{-4}$ for the transport network ($\mathcal{T}_{\phi}$) and $2\times 10^{-4}$ for the potential network ($\varphi_{\theta}$). The inner iteration parameter $N_{c}$ was consistently set to 1. A learning rate decay schedule applied a factor of 10 reduction after each block of 30 training epochs. The underlying architecture for $\mathcal{T}_{\phi}$ was adapted based on the dataset: using the E2E-VarNet backbone [sriram2020end] for simulated MRI data, the PromptMR backbone [PromptMR] for real prospective MRI data, and the MC-CDic backbone [lei2023deep] for clinical prospective CT data. Experiments were performed using PyTorch on an NVIDIA 4090 GPU. The source code will be publicly released following possible publication.
@@
In MRI reconstruction experiments, we conducted comparisons against several existing methods, including the traditional CS-wavelet [lustig2007sparse], two representative deep learning unfolding networks, E2E-VarNet [sriram2020end] and PromptMR [PromptMR], a diffusion-based method (DDS [chung2024decomposed]) and three comparative methods based on OT: OT-CycleGAN [sim2020optimal], UAR [mukherjee2021end] and FSBM [theodoropoulos2025feedback].
For CT reconstruction, our set of comparison methods consists of the traditional BM3D [dabov2007image], two representative deep learning methods (ISTA-Net [lei2023deep], an unfolding network, and REDCNN [chen2017low]), and three OT-based approaches, OT-CycleGAN [sim2020optimal], UAR [mukherjee2021end] and FSBM [theodoropoulos2025feedback].
Motivation for Transport Network Implementation
Our motivation for the implementation of the transport network, $\mathcal{T}_{\phi}$, is guided by three core principles: leveraging task-specific knowledge, ensuring fair comparisons, and demonstrating the broad applicability of our KIDOT framework.
First, we utilize backbones incorporated with task-specific knowledge. In medical image reconstruction, different imaging modalities, such as MRI and CT, possess unique physical properties and data characteristics. For example, MRI heavily relies on multi-coil acquisition, where effectively leveraging the physical information from coil sensitivity maps is often critical for optimal reconstruction. Meanwhile, the imaging mechanism in CT and MRI are different. CT images are produced by Radon Transform and MR images are produced by Fourier Transform. This necessity to incorporate modality-specific knowledge has led the research community to develop specialized, state-of-the-art architectures for each task. To evaluating the performance of our approach for each task, we adopt the state-of-the-art architectures for each task. Note that, for MRI, E2E-VarNet is more light-weight than PromptMR, and we use E2E-VarNet for simulated MRI for the sake of efficient computation.
Second, we ensure a fair comparison in each experiment. For any given task, the transport network in our KIDOT framework uses the same backbone architecture as the corresponding supervised baseline it is compared against. For example, when comparing with E2E-VarNet on the simulated MRI dataset, our KIDOT model also uses an E2E-VarNet-based architecture for its transport network $\mathcal{T}_{\phi}$.
Finally, this approach allows us to demonstrate the versatile usage of KIDOT on different backbones. By deliberately choosing backbones that are identical or highly similar to the supervised baselines, we effectively isolate the contribution of our framework. This ensures that the observed performance gain is attributable to our novel knowledge-informed dynamic optimal transport formulation itself, rather than simply employing a more powerful (and different) network architecture.
@@ B.2 Datasets Details
[…]
@@
Given the potential for anatomical misalignment in real prospective acquisitions (due to different scan directions/timings), we also generated perfectly-paired Simulated Low-Dose CT (SLDCT) images by adding noise in the projection domain. We inject compound Gaussian and Poisson noise into the uncorrupted projections by using the method in [zeng2015simple, li2023noise]. The incident scale photon flux is 40000.We adopt a fan-beam geometry to simulate the uncorrupted projections from the CT images. The geometrical parameters for projection were: source-to-isocenter distance 570.0 mm, source-to-detector distance 962.9 mm, image size 512 × 512 (with pixel spacing 0.6934 × 0.6934 $\text{mm}^{2}$), detector bin number 864, detector bin width 1.0336 mm, and 1200 projection views acquired over a 360-degree orbit.
@@ B.3 Additional Ablation Studies
B.3.1 Effect of Different Number of Discretization Steps NN
We conduct an ablation study to investigate the sensitivity of our method to the number of discretization steps $N$, which is a key hyperparameter in our framework. The choice of $N$ represents a critical trade-off between the fidelity of the ODE approximation and the associated computational cost. A smaller $N$ can lead to a coarse and inaccurate discretization of the continuous transport path, resulting in suboptimal reconstruction quality. Conversely, a large $N$ increases the computational burden, as inference time and memory grow linearly with $N$, and can also introduce training instability due to the risk of gradient exploding or vanishing in deeper backpropagation paths.
[…]
6 35.31 2.01
9 30.06 1.62
12 28.26 1.12
15 29.23 1.31
B.3.2 Sensitivity to the Supervised Trade-off Parameter $\gamma$
[…]
@@ B.3.2 Sensitivity to the Supervised Trade-off Parameter $\gamma$
B.3.3 Effect of Different Lagrange Multiplier $\lambda$
[…]
@@ B.3.3 Effect of Different Lagrange Multiplier $\lambda$
B.3.4 Effect of the Physical Forward Operator $\mathcal{A}$ in Transport Equation
[…]
@@ B.4 Statistical Significance Analysis
[…]
@@
[…]
CS-wavelet <0.001<0.001 <0.001<0.001 <0.001<0.001 <0.001<0.001
E2E-VarNet <0.001<0.001 <0.001<0.001 <0.001<0.001 <0.001<0.001
OT-CycleGAN <0.001<0.001 <0.001<0.001 <0.001<0.001 <0.001<0.001
UAR $0.024$ <0.001<0.001 <0.001<0.001 $0.009$
B.5 Training Cost Curves for Three Datasets
In Figure 7, we display the cost curves over three datasets (i.e., simulated MRI data, real prospective MRI data and clinical prospective CT data) of $\mathcal{T}_{\phi}$ and $\varphi_{\theta}$ in the training process. The $\mathcal{T}_{\phi}$ cost curve is normalized to [0, 1]. $\varphi_{\theta}$ cost is scaled to [0, 1] and then take the negative.
@@ B.5 Training Cost Curves for Three Datasets
B.6 Performance in Low-data Regimes
To evaluate the robustness of our method under data scarcity, we conduct an additional ablation study in a low-data regime. We intentionally restrict the amount of available training data by using reduced subsets of the simulated MRI dataset. Specifically, we train our model and the baselines on only 20% and 10% of the original paired data.
The results are presented in Table 10. In both settings, our KIDOT framework consistently outperforms the E2E-VarNet and UAR baselines across all metrics. This improved performance under data scarcity can be directly attributed to the regularization imposed by KIDOT’s physics-informed constraints. Unlike purely data-driven methods that are prone to overfitting on small datasets, our framework’s adherence to the physical model provides a robust inductive bias, significantly reducing its dependency on the volume of training data and enhancing its generalization capability.
Table 10: Performance comparison in low-data regimes on the Simulated MRI data. We train all methods using 20% and 10% of the original training paired data. Best results for each setting are bolded.
[…]
20% E2E-VarNet 30.97 0.7978 72.43 4.02
UAR 31.30 0.8054 40.32 2.95
KIDOT (Ours) 31.75 0.8131 31.61 2.32
10% E2E-VarNet 29.92 0.7681 92.41 6.07
UAR 30.64 0.7857 87.07 5.68
KIDOT (Ours) 30.88 0.7952 73.14 4.11
B.7 Generalization to More Aggressive Out-of-Distribution (OOD) or Heterogeneous Data
To further investigate the robustness and generalization capabilities of our model, we conducted two challenging cross-domain experiments on the fastMRI dataset, evaluating performance on both cross-anatomy and cross-contrast tasks. These experiments are designed to assess how well the models learn the fundamental principles of MRI reconstruction, rather than simply memorizing features specific to the training data distribution.
Cross-Anatomy (Knee $\rightarrow$ Brain).
In this setup, the models are tasked with reconstructing brain images after being trained on knee data. For the supervised baseline (E2E-VarNet), we used paired knee scans for training. For our unpaired method, KIDOT, we trained it using unpaired data consisting of high-quality knee scans and low-quality brain scans. Both models were then evaluated on their ability to reconstruct brain scans. As shown in Table 11, KIDOT significantly outperforms E2E-VarNet, demonstrating its ability to generalize anatomical knowledge.
Table 11: Generalization performance on the Cross-Anatomy (Knee $\rightarrow$ Brain) task. Best results are bolded.
[…]
E2E-VarNet 33.82 0.7880 80.21 4.24
KIDOT (Ours) 34.62 0.8145 26.87 1.53
Cross-Contrast (T1 $\rightarrow$ T2).
Similarly, this experiment tests generalization across different MRI contrasts. The models were trained on a dataset of unpaired high-quality T1-weighted brain scans and low-quality T2-weighted brain scans. The evaluation was then performed on the task of reconstructing T2-weighted brain scans. The results are presented in Table 12. Again, KIDOT shows superior performance, indicating its robustness to variations in image contrast.
Table 12: Generalization performance on the Cross-Contrast (T1 $\rightarrow$ T2) task. Best results are bolded.
[…]
E2E-VarNet 32.06 0.8229 31.29 1.42
KIDOT (Ours) 32.48 0.8348 29.74 1.33
The superior performance of KIDOT in both of these challenging OOD scenarios provides strong evidence for the benefit of its physics-guided inductive bias. This bias encourages the model to learn a generalizable reconstruction process grounded in the principles of MR imaging, rather than overfitting to the specific anatomical or contrast features of the training domain. While a standard supervised baseline performs well on in-domain data, it struggles when faced with a significant domain shift, as its learned prior is purely data-driven and less generalizable.
B.8 Additional Visual Results
B.8.1 Additional Visual Comparison on Simulated MRI Data
Figure 8 presents additional visual comparisons on simulated MRI data, highlighting that our method reconstructs clearer structural details with fewer artifacts in the regions of interest (red boxes).
@@ B.8.1 Additional Visual Comparison on Simulated MRI Data
B.8.2 Additional Visual Comparison on Real Prospective MRI data
Figures 9 and 10 display more visual results of the compared methods.
@@ B.8.2 Additional Visual Comparison on Real Prospective MRI data
B.8.3 Additional Visual Comparison on Clinical Prospective CT Data
Figures 11 and 12 displays additional visual results of the compared methods on the clinical prospective CT data. As shown in the regions of interest, our method (KIDOT) achieves a better trade-off between noise suppression and structural detail preservation.